Problems (dataset)。 Help Seve&Evit Subm Problem 8-2 to expand the present operation by adding one new outlet She has studied three locations 0 per sandwich Sandwiches sell equipment costs would be $5,.850 per month for ocation A 55,925per monh for locasion B The owner of Genuine Subs, Inc, hopes Each wouid have the same labor and materials costs (food, serving containers, napkins, etc.) c for $3.60 each in all locations. Rent and and $6,175 per month for location C of $2.8 a. Determine the volume necessary at each location to realize a Round your answer to the nearest whole number.) monthly profit of $12,250 (Do not round intermediate calculations. b-1. t expected sales at A, B, and C are 23.250 per month, 26,250 per month,and 25.250 per month, respectively,calculate the prote of the each locations? (Omit the $ sign in your response.) <Prev 2of8田Neg
Answer #1

1.Volume required for profit of $12250 at location A=(cost of location+monthly profit)/(selling price-material cost)

=(5850+12250)÷(3.6-2.8)=22625 units

2.Volume required for profit of $12250 at location B=(cost of location+monthly profit)/(selling price-material cost)

=(5925+12250)÷(3.6-2.8)=22718.75 units

3.Volume required for profit of $12250 at location A=(cost of location+monthly profit)/(selling price-material cost)

=(6175+12250)÷(3.6-2.8)=23031.25 units

B.Profit at shop A =23250×(3.6-2.8)-5850=$12750

Profit at shop B=26250×0.8-5925=$15075

Profit at shop C=25250×0.8-6175=$14025